# Settling Constant Regrets in Linear Markov Decision Processes

### v1

@@ Abstract
We study the constant regret guarantees in reinforcement learning (RL). Our objective is to design an algorithm that incurs only finite regret over infinite episodes with high probability. We introduce an algorithm, Cert-LSVI-UCB, for misspecified linear Markov decision processes (MDPs) where both the transition kernel and the reward function can be approximated by some linear function up to misspecification level $\zeta$. At the core of Cert-LSVI-UCB is an innovative certified estimator, which facilitates a fine-grained concentration analysis for multi-phase value-targeted regression, enabling us to establish an instance-dependent regret bound that is constant w.r.t. the number of episodes. Specifically, we demonstrate that for an MDP characterized by a minimal suboptimality gap $\Delta$, Cert-LSVI-UCB has a cumulative regret of $\widetilde{\mathcal{O}}(d^{3}H^{5}/\Delta)$ with high probability, provided that the misspecification level $\zeta$ is below $\widetilde{\mathcal{O}}(\Delta/(\sqrt{d}H^{2}))$. Remarkably, this regret bound remains constant relative to the number of episodes $K$. To the best of our knowledge, Cert-LSVI-UCB is the first algorithm to achieve a constant, instance-dependent, high-probability regret bound in RL with linear function approximation for infinite runs without relying on prior distribution assumptions. This not only highlights the robustness of Cert-LSVI-UCB to model misspecification but also introduces novel algorithmic designs and analytical techniques of independent interest.
@@ 1 Introduction
Reinforcement learning (RL) has been a popular approach for teaching agents to make decisions based on feedback from the environment. RL has shown great success in a variety of applications, including robotics (Kober et al., 2013), gaming (Mnih et al., 2013), and autonomous driving. In the most of these applications, there is a common expectation that RL agents will master tasks with making only a bounded number of mistakes, even over indefinite runs. However, theoretical support of this expectation is limited in RL theory literature: in the instance-independent case, existing works, to mention a few, Jin et al. (2020); Ayoub et al. (2020); Wang et al. (2019), provided only $\widetilde{\mathcal{O}}(\sqrt{K})$ regret upper bounds; in the instance-dependent setting, Simchowitz and Jamieson (2019); Yang et al. (2021); He et al. (2021a) provided logarithmic $\widetilde{\mathcal{O}}(\Delta^{-1}\log K)$ high-probability regret upper bounds for both tabular MDPs and MDPs with linear function approximations, given a suboptimality gap $\Delta$. However, these findings suggest that an agent’s regret increases with the number of episodes $K$, contradicting the practical expectation of finite mistakes. Conversely, recent years have witnessed a series of work providing a constant regret bound for RL and bandits, suggesting that an RL agent’s regret may remain bounded even when it faces an indefinite number of episodes. Papini et al. (2021a); Zhang et al. (2021) have provided instance-dependent constant regret bound under the assumption of prior data distribution. However, verifying these data distribution assumptions can be difficult or infeasible. On the other hand, it is known that high-probability constant regret bound can be achieved unconditionally in multi-armed bandits (Abbasi-Yadkori et al., 2011) and in contextual linear bandits if and only if the misspecification is sufficiently small with respect to the minimal sub-optimality gap (Zhang et al., 2023b). This raises a critical question:
@@
To answer this question, we introduce a novel algorithm, which we refer to as Cert-LSVI-UCB, for reinforcement learning with linear function approximation. To encompass a broader range of real-world scenarios characterized by large state-action spaces and the need for function approximations, we adapt the misspecified linear MDP (Jin et al., 2020) setting, where both the transition kernel and reward function can be approximated by a linear function with approximation error $\zeta$. We show that, with our innovative design of certified estimator and novel analytical techniques, Cert-LSVI-UCB achieves constant regret without relying on any prior assumption on data distributions. Our key contributions are summarized as follows:
  • We introduce a parameter-free algorithm, referred to as Cert-LSVI-UCB, featuring a novel certified estimator for testing when the confidence set fails. This innovative approach enables Cert-LSVI-UCB to achieve a constant, instance-dependent, high probability regret bound of $\widetilde{\mathcal{O}}(d^{3}H^{5}/\Delta)$ for tasks with a suboptimality gap $\Delta$, under the condition that the misspecification level $\zeta$ is bounded by $\zeta<\widetilde{\mathcal{O}}\big{(}\Delta/(\sqrt{d}H^{2})\big{)}$. This bound is termed a high probability constant regret bound, indicating that it does not depend on the number of episodes $K$. We note that this constant regret bound matches the logarithmic expected regret lower bound of $\Omega(\Delta^{-1}\log K)$, suggesting that our result is valid and optimal in terms of the dependence of suboptimality gap $\Delta$.
@@
Notation. Scalars are denoted by lower case letter. Vectors are denoted by lower case boldface letters $\mathbf{x}$, and matrices by upper case boldface letters $\mathbf{A}$. We denote by $[k]$ the set $\{1,2,\cdots,k\}$ for positive integers $k$. We use $\log x$ to denote the logarithm of $x$ to base $2$. For two non-negative sequence $\{a_{n}\},\{b_{n}\}$, $a_{n}\leq\mathcal{O}(b_{n})$ means that there exists a positive constant $C$ such that $a_{n}\leq Cb_{n}$; $a_{n}\leq\widetilde{\mathcal{O}}(b_{n})$ means there exists a positive constant $k$ such that $a_{n}\leq\mathcal{O}(b_{n}\log^{k}b_{n})$; $a_{n}\geq\Omega(b_{n})$ means that there exists a positive constant $C$ such that $a_{n}\geq Cb_{n}$; $a_{n}\geq\widetilde{\Omega}(b_{n})$ means there exists a positive constant $k$ such that $a_{n}\geq\Omega(b_{n}\log^{-k}b_{n})$; $a_{n}\geq\omega(b_{n})$ means that $\lim_{n\rightarrow\infty}b_{n}/a_{n}=0$. For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and a positive semi-definite matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$, we define $\|\mathbf{x}\|_{\mathbf{A}}^{2}=\mathbf{x}^{\top}\mathbf{A}\mathbf{x}$. For any set $\mathcal{C}$, we use $|\mathcal{C}|$ to denote its cardinality. We denote the identity matrix by $\mathbf{I}$ and the empty set by $\emptyset$. The total variation distance of two distribution measures $\mathbb{P}(\cdot)$ and $\mathbb{Q}(\cdot)$ is denoted by $\|\mathbb{P}(\cdot)-\mathbb{Q}(\cdot)\|_{\mathrm{TV}}$.
@@ 2 Related Work
Algorithm Misspecified MDP? Result
@@
Table 1: Instance-dependent regret bounds for different algorithms under the linear MDP setting. Here $d$ is the dimension of the linear function $\bm{\phi}(s,a)$, $H$ is the horizon length, $\Delta$ is the minimal suboptimality gap. All results in the table represent high probability regret bounds. The regret bound depends the number of episodes $K$ in He et al. (2021a) and the minimum positive eigenvalue $\lambda$ of features mapping in Papini et al. (2021b). Misspecified MDP? indicates if the algorithm can (✓) handle the misspecified linear MDP or not ($\times$).
@@ 3 Preliminaries
For a $\zeta$-approximate linear MDP, for any policy $\pi$, there exist corresponding weights $\{\mathbf{w}_{h}^{\pi}\}_{h\in[H]}$ where $\mathbf{w}_{h}^{\pi}=\bm{\theta}_{h}+\int V_{h+1}^{\pi}(s^{\prime})\mathrm{d}\bm{\mu}_{h}(s^{\prime})$ such that for any $(s,a,h)\in{\mathcal{S}}\times\mathcal{A}\times[H]$, $\big{|}Q_{h}^{\pi}(s,a)-\left\langle\bm{\phi}(s,a),\mathbf{w}_{h}^{\pi}\right\rangle\big{|}\leq 2H\zeta$. We have $\|wb_{h}^{\pi}\|_{2}\leq 2H\sqrt{d}$.
@@ 4.1 Main algorithm: Cert-LSVI-UCB
1:  Set $V^{k}_{H+1}(s)=0$ for all $(s,k)\in{\mathcal{S}}\times[K]$, $\mathcal{C}^{k}_{h,l}=\emptyset$ for all $(h,l)\in[H]\times\mathbb{N}^{+},\lambda=16$
@@
${\mathbf{w}}^{k}_{h,l}\leftarrow\mathop{\mathrm{argmin}}_{\mathbf{w}\in\mathbb{R}^{d}}\lambda\|\mathbf{w}\|_{2}^{2}+\sum_{\tau\in\mathcal{C}_{h,l}^{k-1}}\big{(}\mathbf{w}^{\top}\bm{\phi}_{h}^{\tau}-r_{h}^{\tau}-\widehat{V}_{h+1}^{k}(s_{h+1}^{\tau})\big{)}^{2}.$
@@ 4.2 Subroutine: Cert-LinUCB
11:     else if $\max\big{\{}V^{k}_{h,l}(s)-3\cdot 2^{-l},\widecheck{V}^{k}_{h,l-1}(s)\big{\}}>\min\big{\{}V^{k}_{h,l}(s)+3\cdot 2^{-l},\widehat{V}^{k}_{h,l-1}(s)\big{\}}$ then
@@
  • Condition 3: In Line 11, we compare the value of the pessimistic value function $\widecheck{V}_{h,l}^{k}(s)$ and the optimistic value function $\widehat{V}_{h,l}^{k}(s)$ which will be assigned in Line 14 and Line 15, if the pessimistic estimation will be greater than the optimistic estimation, we will stop at that phase and take the greedy policy on previous phase $\pi_{h}^{k}(s)=\pi_{h,l-1}^{k}(s)$. Only in this case, the Algorithm 2 outputs flag $f_{h}^{k}(s)=0$, which means this observation will not be used in Line 15 in Algorithm 1.
@@
Specially, our algorithm differs from that in Vial et al. (2022) in terms of Condition 3 to certify the performance of the estimation. In particular, a well-behaved estimation should always guarantee that the optimistic estimation is greater than the pessimistic estimation. According to Line 14 and Line 15, this is equivalent to the confidence region for $l$-th phase has intersection of the previous confidence region $[\widecheck{V}_{h,l-1}^{k}(s),\widehat{V}_{h,l-1}^{k}(s)]$. Otherwise, we hypothesis the estimation on $l$-th phase is corrupted by either misspecification or bad concentration event, thus will stop the algorithm. We will revisit the detail of this design later.
@@ 5 Constant Regret Guarantee
Under Assumption 3.1, let $\gamma_{l}=5(l+20+\lceil\log(ld)\rceil)dH\sqrt{\log(16ldH/\delta)}$ for some fixed $40<\delta<1/$ 4. With probability at least $1-4\delta$, if the minimal suboptimality gap $\Delta$ satisfies $\Delta>\widetilde{\Omega}\big{(}\sqrt{d}H^{2}\zeta\big{)}$, then for all $K\in\mathbb{N}^{+}$, the regret of Algorithm 1 is upper bounded by
@@
6 Highlight of Proof Techniques
@@
6.1 Technical challenges
Challenge 1. Achieving layer-wise local estimation error.
In the analysis of the value function under misspecified linear MDPs, we need to follow the multi-phase estimation strategy (Vial et al., 2022) to eliminate suboptimal actions and improve the robustness of the next phase estimation. Similar approaches have been observed in Zhang et al. (2023b); Chu et al. (2011) within the framework of (misspecified) linear bandits. However, unlike linear bandits, when constructing the empirical value function $\widehat{V}_{h}$ for stage $h$ in linear MDPs, Jin et al. (2020) requires a covering statement on value functions to ensure the convergence of the regression, which is written by: (see Lemma D.4 in Jin et al. (2020) for details)
@@ 6.1 Technical challenges
where we employ notation $\widetilde{\mathcal{O}}_{H}$ to obscure the dependence on $H$ to simplify the presentation. We use the notation $\mathcal{V}_{h+1}^{k}$ to denote as an $\kappa$-covering for the value functions $\widehat{V}_{h+1}^{k}$. Takemura et al. (2021); Vial et al. (2022) used quantification instead of the covering number, but this approach still encounters the issue of taking the union bound across the set of value functions, thereby incorporating the dependency on the cardinality of this set.
In the multi-phase algorithm, the regression employs a distinct set $\mathcal{C}=\mathcal{C}_{h,l}^{k}$ for each phase $l$. However, all these regressions use the overall empirical value function $\widehat{V}_{h+1}^{k}$ from the subsequent stage $h+1$, which is formulated using all pairs of parameters $\big{\{}\mathbf{w}_{h,\ell}^{k},\mathbf{U}_{h,\ell}^{k}\big{\}}_{\ell}$ in $L$ phases. Consequently, the covering number $\log|\mathcal{V}_{h+1}^{k}|$ is directly proportional to the number of phases $L=\mathcal{O}(\log K)$.
Therefore, when analyzing any single phase $l$, prior analysis cannot eliminate the $\log K$ term from (6.1) to achieve a local estimation error independent that is independent of the logarithmic number of global episodes $\log K$. Furthermore, due to the algorithm design of previous methods (Vial et al., 2022), additional $\log K$ terms may be introduced, induced by global quantification (i.e., $\varepsilon_{tol}=d/\sqrt{K})$.
Challenge 2. Achieving constant regret from local estimation error

### v2

# Achieving Constant Regret in Linear Markov Decision Processes

@@ Abstract
We study the constant regret guarantees in reinforcement learning (RL). Our objective is to design an algorithm that incurs only finite regret over infinite episodes with high probability. We introduce an algorithm, Cert-LSVI-UCB, for misspecified linear Markov decision processes (MDPs) where both the transition kernel and the reward function can be approximated by some linear function up to misspecification level $\zeta$. At the core of Cert-LSVI-UCB is an innovative certified estimator, which facilitates a fine-grained concentration analysis for multi-phase value-targeted regression, enabling us to establish an instance-dependent regret bound that is constant w.r.t. the number of episodes. Specifically, we demonstrate that for a linear MDP characterized by a minimal suboptimality gap $\Delta$, Cert-LSVI-UCB has a cumulative regret of $\widetilde{\mathcal{O}}(d^{3}H^{5}/\Delta)$ with high probability, provided that the misspecification level $\zeta$ is below $\widetilde{\mathcal{O}}(\Delta/(\sqrt{d}H^{2}))$. Here $d$ is the dimension of the feature space and $H$ is the horizon. Remarkably, this regret bound is independent of the number of episodes $K$. To the best of our knowledge, Cert-LSVI-UCB is the first algorithm to achieve a constant, instance-dependent, high-probability regret bound in RL with linear function approximation without relying on prior distribution assumptions.
@@ 1 Introduction
Reinforcement learning (RL) has been a popular approach for teaching agents to make decisions based on feedback from the environment. RL has shown great success in a variety of applications, including robotics (Kober et al., 2013), gaming (Mnih et al., 2013), and autonomous driving. In most of these applications, there is a common expectation that RL agents will master tasks after making only a bounded number of mistakes, even over indefinite runs. However, theoretical support for this expectation is limited in RL literature: in the worst case, existing works such as Jin et al. (2020); Ayoub et al. (2020); Wang et al. (2019) only provided $\widetilde{\mathcal{O}}(\sqrt{K})$ regret upper bounds with $K$ being the number of episodes; in the instance-dependent case, Simchowitz and Jamieson (2019); Yang et al. (2021); He et al. (2021a) achieved logarithmic high-probability regret upper bounds (e.g., $\widetilde{\mathcal{O}}(\Delta^{-1}\log K)$) for both tabular MDPs and MDPs with linear function approximations, provided a minimal suboptimality gap $\Delta$. However, these findings suggest that an agent’s regret increases with the number of episodes $K$, contradicting to the expectation of finite mistakes in practice. To close this gap between theory and practice, there is a recent line of work proving constant regrets bound for RL and bandits, suggesting that an RL agent’s regret may remain bounded even when it encounters an indefinite number of episodes. Papini et al. (2021a); Zhang et al. (2021) have provided instance-dependent constant regret bound under certain coverage assumptions on the data distribution. However, verifying these data distribution assumptions can be difficult or even infeasible. On the other hand, it is known that high-probability constant regret bound can be achieved unconditionally in multi-armed bandits (Abbasi-Yadkori et al., 2011) and contextual linear bandits if and only if the misspecification is sufficiently small with respect to the minimal sub-optimality gap (Zhang et al., 2023b). This raises a critical question:
@@
To answer this question, we introduce a novel algorithm, which we refer to as Cert-LSVI-UCB, for reinforcement learning with linear function approximation. To encompass a broader range of real-world scenarios characterized by large state-action spaces and the need for function approximation, we consider the misspecified linear MDP (Jin et al., 2020) setting, where both the transition kernel and reward function can be approximated by a linear function with approximation error $\zeta$. We show that, with our innovative design of certified estimator and novel analysis, Cert-LSVI-UCB achieves constant regret without relying on any prior assumption on data distributions. Our key contributions are summarized as follows:
  • We introduce a parameter-free algorithm, referred to as Cert-LSVI-UCB, featuring a novel certified estimator for testing when the confidence set fails. This certified estimator enables Cert-LSVI-UCB to achieve a constant, instance-dependent, high probability regret bound of $\widetilde{\mathcal{O}}(d^{3}H^{5}/\Delta)$ for tasks with a suboptimality gap $\Delta$, under the condition that the misspecification level $\zeta$ is bounded by $\zeta<\widetilde{\mathcal{O}}\big{(}\Delta/(\sqrt{d}H^{2})\big{)}$. This bound is termed a high probability constant regret bound, indicating that it does not depend on the number of episodes $K$. We note that this constant regret bound matches the logarithmic expected regret lower bound of $\Omega(\Delta^{-1}\log K)$, suggesting that our result is valid and optimal in terms of the dependence on the suboptimality gap $\Delta$.
@@
Notation. Vectors are denoted by lower case boldface letters such as $\mathbf{x}$, and matrices by upper case boldface letters such as $\mathbf{A}$. We denote by $[k]$ the set $\{1,2,\cdots,k\}$ for positive integers $k$. We use $\log x$ to denote the logarithm of $x$ to base $2$. For two non-negative sequence $\{a_{n}\},\{b_{n}\}$, $a_{n}\leq\mathcal{O}(b_{n})$ means that there exists a positive constant $C$ such that $a_{n}\leq Cb_{n}$; $a_{n}\leq\widetilde{\mathcal{O}}(b_{n})$ means there exists a positive constant $k$ such that $a_{n}\leq\mathcal{O}(b_{n}\log^{k}b_{n})$; $a_{n}\geq\Omega(b_{n})$ means that there exists a positive constant $C$ such that $a_{n}\geq Cb_{n}$; $a_{n}\geq\widetilde{\Omega}(b_{n})$ means there exists a positive constant $k$ such that $a_{n}\geq\Omega(b_{n}\log^{-k}b_{n})$; $a_{n}\geq\omega(b_{n})$ means that $\lim_{n\rightarrow\infty}b_{n}/a_{n}=0$. For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and a positive semi-definite matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$, we define $\|\mathbf{x}\|_{\mathbf{A}}^{2}=\mathbf{x}^{\top}\mathbf{A}\mathbf{x}$. For any set $\mathcal{C}$, we use $|\mathcal{C}|$ to denote its cardinality. We denote the identity matrix by $\mathbf{I}$ and the empty set by $\varnothing$. The total variation distance of two distribution measures $\mathbb{P}(\cdot)$ and $\mathbb{Q}(\cdot)$ is denoted by $\|\mathbb{P}(\cdot)-\mathbb{Q}(\cdot)\|_{\mathrm{TV}}$.
@@ 2 Related Work
Algorithm Allow Misspecification? Result
@@
Table 1: Instance-dependent regret bounds for different algorithms under the linear MDP setting. Here $d$ is the dimension of the linear function $\bm{\phi}(s,a)$, $H$ is the horizon length, $\Delta$ is the minimal suboptimality gap. All results in the table represent high probability regret bounds. The regret bound depends the number of episodes $K$ in He et al. (2021a) and the minimum positive eigenvalue $\lambda$ of features mapping in Papini et al. (2021b). Allow Misspecification? indicates if the algorithm can (✓) handle the misspecified linear MDP or not ($\times$).
@@ 3 Preliminaries
For a $\zeta$-approximate linear MDP, for any policy $\pi$, there exist corresponding weights $\{\mathbf{w}_{h}^{\pi}\}_{h\in[H]}$ where $\mathbf{w}_{h}^{\pi}=\bm{\theta}_{h}+\int V_{h+1}^{\pi}(s^{\prime})\mathrm{d}\bm{\mu}_{h}(s^{\prime})$ such that for any $(s,a,h)\in{\mathcal{S}}\times\mathcal{A}\times[H]$, $\big{|}Q_{h}^{\pi}(s,a)-\left\langle\bm{\phi}(s,a),\mathbf{w}_{h}^{\pi}\right\rangle\big{|}\leq 2H\zeta$. We have $\|\mathbf{w}_{h}^{\pi}\|_{2}\leq 2H\sqrt{d}$.
@@ 4.1 Main algorithm: Cert-LSVI-UCB
1:  Set $V^{k}_{H+1}(s)=0$ for all $(s,k)\in{\mathcal{S}}\times[K]$, $\mathcal{C}^{k}_{h,l}=\varnothing$ for all $(h,l)\in[H]\times\mathbb{N}^{+},\lambda=16$
@@
$\textstyle{{\mathbf{w}}^{k}_{h,l}\leftarrow\mathop{\mathrm{argmin}}_{\mathbf{w}\in\mathbb{R}^{d}}\lambda\|\mathbf{w}\|_{2}^{2}+\sum_{\tau\in\mathcal{C}_{h,l}^{k-1}}\big{(}\mathbf{w}^{\top}\bm{\phi}_{h}^{\tau}-r_{h}^{\tau}-\widehat{V}_{h+1}^{k}(s_{h+1}^{\tau})\big{)}^{2}.}$
@@ 4.2 Subroutine: Cert-LinUCB
[…]
@@
[…]
@@
[…]
@@ 5 Constant Regret Guarantee
Under Assumption 3.1, let $\gamma_{l}=5(l+20+\lceil\log(ld)\rceil)dH\sqrt{\log(16ldH/\delta)}$ for some fixed $40<\delta<1/$ 4. With probability at least $1-4\delta$, if misspecification level $\zeta$ is below $\widetilde{\mathcal{O}}\big{(}\Delta/(\sqrt{d}H^{2})\big{)}$ where $\Delta$ is the minimal suboptimality gap, then for all $K\in\mathbb{N}^{+}$, the regret of Algorithm 1 is upper bounded by
@@
6 Technical Challenges and Highlight of Proof Techniques
@@
6.1 Challenge 1. Achieving layer-wise local estimation error.
In the analysis of the value function under misspecified linear MDPs, we follow the multi-phase estimation strategy (Vial et al., 2022) to eliminate suboptimal actions and improve the robustness of the next phase estimation. Similar approaches have been observed in Zhang et al. (2023b); Chu et al. (2011) within the framework of (misspecified) linear bandits. However, unlike linear bandits, when constructing the empirical value function $\widehat{V}_{h}$ for stage $h$ in linear MDPs, Jin et al. (2020) requires a covering statement on value functions to ensure the convergence of the regression, which is written by: (see Lemma D.4 in Jin et al. (2020) for details)
@@
where we employ notation $\widetilde{\mathcal{O}}_{H}$ to obscure the dependence on $H$ to simplify the presentation. We use the notation $\mathcal{V}_{h+1}^{k}$ to denote as an $\kappa$-covering, (or quantification in Takemura et al. (2021); Vial et al. (2022)) for the value functions $\widehat{V}_{h+1}^{k}$. However, in the multi-phase algorithm, the empirical value function $\widehat{V}_{h+1}^{k}$ from the subsequent stage $h+1$, which is formulated using all pairs of parameters $\big{\{}\mathbf{w}_{h,\ell}^{k},\mathbf{U}_{h,\ell}^{k}\big{\}}_{\ell}$ in $L$ phases. Consequently, the covering number $\log|\mathcal{V}_{h+1}^{k}|$ is directly proportional to the number of phases $L=\mathcal{O}(\log K)$.
Therefore, when analyzing any single phase $l$, prior analysis cannot eliminate the $\log K$ term from (6.1) to achieve a local estimation error independent that is independent of the logarithmic number of global episodes $\log K$. Furthermore, due to the algorithm design of previous methods (Vial et al., 2022), additional $\log K$ terms may be introduced by global quantification (i.e., $\varepsilon_{tol}=d/\sqrt{K})$.
Our approach: Cert-LinUCB.
To tackle this challenge, we introduce the certified estimator into Algorithm 2 and use a ‘local quantification’ to ensure the quantification error of each phase $l$ depend on the local phase $\widetilde{\mathcal{O}}(l)$ instead of the global parameter $\log K$. The certified estimator works as follows: Considering the concentration term we need to control for each phase $l$:
$\Big{\|}\textstyle{\sum_{\tau\in\mathcal{C}_{h,l}^{k}}}\bm{\phi}^{\tau}_{h}\big{[}\widehat{V}_{h+1}^{k}(s^{\tau})-\mathbb{E}[\widehat{V}_{h+1}^{k}(s^{\tau})]\big{]}\Big{\|}_{(\mathbf{U}^{k}_{h,l})^{-1}},$ (6.2)
Lemma 6.1 (Lemma C.4, informal).
Let $\widehat{V}_{h+1,l_{+}}^{k}$ be the output of Algorithm 2 terminated at phase $l_{+}\in\mathbb{N}^{+}$, then with probability is at least $1-2\delta$,
Lemma 6.1 suggests a concentration inequality at any phase $l_{+}$, and the following lemma suggests that this procedure will only introduce an $\widetilde{\mathcal{O}}(2^{-l_{+}})$ error, under some faithful extension of the $\widehat{V}^{k}_{h,l_{+}}(s)$:
Lemma 6.2 (Lemma C.2, informal).
For any $l_{+}\in\mathbb{N}^{+}$, $|\widehat{V}^{k}_{h}(s)-\widehat{V}^{k}_{h,l_{+}}(s)|\leq 6\cdot 2^{-l_{+}}$.
Therefore, if a large enough $l_{+}$ can be reached in Algorithm 2, combining Lemma 6.1 and Lemma 6.2 allow us to bound (6.2) without introducing $\log K$ factors. The next lemma shows that the Line 11 will only never be triggered in shallow layer $l$.
Lemma 6.3 (Lemma C.8, informal).
Lemma 6.3 delivers a clear message: In the well-specified setting, Line 11 will never be triggered ($l\geq\infty$). When the misspecification level is large, then Line 11 will be more likely triggered, indicating it’s harder for the algorithm to proceed to deeper layer. The contribution of the certified estimator yields the following important lemma regarding the ‘local estimation error’:
Lemma 6.4 (Lemma C.12, Informal).
Remark 6.5.
He et al. (2021b) achieved a similar $\mathcal{O}\big{(}d^{3}H^{5}\varepsilon^{-2}\big{)}$ uniform-PAC bound for (well-specified) linear MDP. Comparing with Lemma 6.4 with $\zeta=0$, one can find that our result is better than He et al. (2021b). In addition, Lemma 6.4 ensures this uniform-PAC result under all stage $h\in[H]$ while He et al. (2021b) only ensure the $h=1$. This improvement is achieved by a more efficient data selection strategy which we will discuss in detail in Appendix B.1.
6.2 Challenge 2. Achieving constant regret from local estimation error